A cargo truck blow its horn and can be heard at a distance of 10.0 m with a sound intensity level 93.0 dB. If 6 identical cargo trucks simultaneously blow their horns, what combined sound intensity level can be heard at a distance of 10.0 m form these trucks?


1
Expert's answer
2022-05-10T13:20:52-0400
"L_1=10\lg\frac{I_1}{I_0}"

"L_6=10\lg\frac{I_6}{I_0}=10\lg\frac{6I_1}{I_0}=10(\lg 6+\lg\frac{I_1}{I_0})"

"L_6=10\lg 6+L_1=10\lg 6+93.0\rm\: dB=101\; dB"